# Algorithmic Collusion at Test Time: A Meta-game Design and Evaluation

### v1

@@ B.2 Detailed Results of UCB
Figure 20: The best-response graph for UCBs. $c_{1}=c_{2}=1$, $t=10,000$.
@@
Table 13: The average payoffs over 40,000 runs per strategy pair of UCB with symmetric cost $c_{1}=c_{2}=1$ and time horizon $t=10,000$. For reference, the competitive payoff is $\bar{r}^{N}=0.22$, and the monopoly payoff is $\bar{r}^{M}=0.34$.
@@
Figure 21: The change in CR ($\Delta_{t}$(CR)) and PC ($\Delta_{t}$(PC)) over time for UCB.
@@
Figure 22: The best-response graph of UCB with the addition of a Q-learning agent (RD, 0.5).
@@ C.3 Detailed Results of LLM
Table 15: The average payoffs over 40 runs per strategy pair of LLMs ($c_{1}=c_{2}=1$ and $t=50$). For reference, in this setting, the competitive and monopoly payoffs are $\bar{r}^{N}=0.22$ and $\bar{r}^{M}=0.34$.
@@
Figure 25: The best-response graph for LLM players. $c_{1}=c_{2}=1$ and payoffs are evaluated at $t=50.$

### v2

@@ B.2 Detailed Results of UCB
Figure 20: Best-response graph for UCBs. $c_{1}=c_{2}=1$, $t=10,000$.
@@
Table 13: Average payoffs over 4,000 runs in the payoff matrix of UCB-based strategies, with symmetric cost $c_{1}=c_{2}=1$ and time horizon $t=10,000$. For reference, the competitive payoff is $\bar{r}^{N}=0.22$, and the monopoly payoff is $\bar{r}^{M}=0.34$.
@@
Figure 21: Changes in CR ($\Delta_{t}$(CR)) and PC ($\Delta_{t}$(PC)) over time for UCB.
@@
Figure 22: Best-response graph of UCB with the addition of a Q-learning agent (RD, 0.5).
@@ C.3 Detailed Results of LLM
Table 15: Average payoffs over 40 runs in the payoff matrix of LLMs ($c_{1}=c_{2}=1$ and $t=50$). For reference, in this setting, the competitive and monopoly payoffs are $\bar{r}^{N}=0.22$ and $\bar{r}^{M}=0.34$.
@@
Figure 25: Best-response graph for LLM players. $c_{1}=c_{2}=1$ and payoffs are evaluated at $t=50.$